OpenStudy (anonymous):

\[f(x)=2(x-2)^{2}-1\] x-intercept Help!!!

OpenStudy (anonymous):

i coulda sworn i did this problem, but lets do it again two ways.

OpenStudy (anonymous):

first way is to expand and get \[2(x-2)(x-2)-1=2(x^2-4x+4)-1=2x^2-8x+8-1\] \[=2x^2-8x+7\]

OpenStudy (anonymous):

now you can use the quadratic formula if you like, or you can complete the square

OpenStudy (anonymous):

x-intercept is where the graph crosses the x-axis. There are two of these, they are called the zero's of x. One is 1.2928932, and the other is: x= 2.707

OpenStudy (anonymous):

do you want to complete the square? it may be easier in this case.

OpenStudy (anonymous):

ok so the x-intercept is 1.3 and 2.7 what would y-intercept be???

OpenStudy (anonymous):

actually the x - intercepts are \[\frac{4\pm\sqrt{2}}{2}\]